OpenStudy (anonymous):

What is the approximate volume of the cone? Use 3.14 for π. A. 2170 cm3 B. 6510 cm3 C. 1206 cm3 D. 3260 cm3

OpenStudy (anonymous):

will give medal @StudyShark2000

OpenStudy (campbell_st):

well the volume of a cone is \[V = \frac{1}{3} \pi r^2h\] so using 3.14 \[V = \frac{1}{3} \times 3.14 \times r^2 \times h\] you need to substitute the radius and height values to find the volume. you should also know the volume of the cone is 1/3 the volume of a cylinder with the same height and radius hope it helps

OpenStudy (anonymous):